OpenStudy (anonymous):

Please Help!!!!! will Fan And Give Medals!!!!!! Calculate the discriminant. x2 – 2x + 4 = 0 A. 0 B. –12 C. 12 D. 20

OpenStudy (anonymous):

@paki

OpenStudy (paki):

use b^2-4ac... and solve...

OpenStudy (anonymous):

I dont know how.

OpenStudy (paki):

here a= 1... b= -2... and c= 4.... now put it in the formula...and solve... what you got...?

OpenStudy (paki):

@sammy34daisy

OpenStudy (acxbox22):

(-2)^2-4(1)(4)= 4-16=??

OpenStudy (anonymous):

-12

OpenStudy (paki):

yup -12

OpenStudy (anonymous):

can you help me with more please?

OpenStudy (paki):

yup, will try... :-)

OpenStudy (anonymous):

What number would you have to add to both sides to complete the square? x2 – 16x = –8 A. –16 B. 16 C. –64 D. 64

OpenStudy (acxbox22):

-16/2=-4 (-4)^2=??

OpenStudy (anonymous):

16

OpenStudy (acxbox22):

yes

OpenStudy (anonymous):

Solve by using the perfect squares method. x2 – 12x + 36 = 0 A. 6 B. 18 C. –18 D. –6

OpenStudy (acxbox22):

(x-6)^2=0 x-6=0 x=??

OpenStudy (anonymous):

-6?

OpenStudy (acxbox22):

close

OpenStudy (anonymous):

A

OpenStudy (acxbox22):

yes

OpenStudy (anonymous):

Solve. x2 – 5x + 6 = 0 A. {3, –2} B. {–2, –3} C. {5, –1} D. {2, 3}